OpenStudy (anonymous):

daisy is shipping a package to herdaughter molly. if the box's dimentions are equal, and the leangth is 3 feet long, what is the volume of the box in cubic feet? a-1 b-3 c-9 d-27

OpenStudy (anonymous):

d lets say the dimensions are Length = x ft Width = x ft height =x ft Then the volume V = Length *Width*Height(ft) V = x^3 But since x = 3 V=3^3 Hence V = 27ft

OpenStudy (anonymous):

Sorry 27 ft^3

OpenStudy (anonymous):

it looks like a bunch of gibberish to me xc im sorry

OpenStudy (anonymous):

ok A box has 3 dimensions They are Length, Width and Height.  If they are equal then we assume each to be assigned the same random variable lets pick the letter Q feet for example Now since Volume is defined as Length * Width * Height, Then Volume = Length * Width * Height, But since we assigned a random letter to the dimensions, Volume = Q * Q * Q = Q^3 But we know that one dimension is 3 ft therefore all dimensions are 3 ft each Hence substituting Q for 3 we get Volume = 3^3 (ft) Volume = 27[ft ^{3}\]